# On polynomial interpolation in the monomial basis

### v4

@@ 2 Polynomial interpolation in the monomial basis
When the Vandermonde system is solved by a backward stable linear system solver, the set of assumptions (27) and (28) is satisfied with constant $\gamma_{N}=\mathcal{O}({\lVert V^{(N)}\rVert}_{2})$, from which it follows that the condition (30) becomes $\kappa(V^{(N)})\lesssim\frac{1}{u}$. Without loss of generality, one can assume that $\Gamma$ is inside the unit disk $D_{1}$ centered at the origin, such that ${\lVert V^{(N)}\rVert}_{2}$ is small. In this case, we observe that $\gamma_{N}\lesssim 1$ for at least $N\leq 100$ when LU factorization with partial pivoting (which is backward stable) is used to solve the Vandermonde system. Therefore, given collocation points with a small Lebesgue constant $\Lambda_{N}$, the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$ is bounded by approximately ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+u{\lVert a^{(N)}\rVert}_{2}$. In Figure 5, we plot the values of ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$, ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}$, and $u{\lVert a^{(N)}\rVert}_{2}$, for functions appear in Section 1, in order to validate the theorem above.
@@
(e) ${\lvert x+0.1\rvert}^{2.5}$
(f) ${\lvert\sin(5x)\rvert}^{3}$
Figure 5: Polynomial interpolation error, monomial approximation error and $u\cdot{\lVert a^{(N)}\rVert}_{2}$. These functions are the ones that appear in Section 1.
@@
The second term on the right-hand side of (32) is an upper bound of the backward error ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$, i.e., the extra loss of accuracy caused by the use of a monomial basis. Note that the absolute condition number of the evaluation of $P_{N}(z)$ in the monomial basis is around ${\lVert a^{(N)}\rVert}_{2}$ when ${\lvert z\rvert}\approx 1$, so that the resulting error is bounded by $u\cdot{\lVert a^{(N)}\rVert}_{2}$, which is always smaller than $2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}$.
The rest of this section is structured as follows. First, we review a classical result on function approximation over a smooth simple arc $\Gamma\subset\mathbbm{C}$ by polynomials. Next, we study the backward error ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$ by bounding the 2-norm of the monomial coefficients of the interpolating polynomial. Finally, we study the growth of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$, which determines the validity of the condition on the a priori error estimate (32).
Below, we define a generalization of the Bernstein ellipse, to the case of a smooth simple arc in the complex plane.
@@
Given a smooth simple arc $\Gamma$ in the complex plane, we define $E_{\rho}$ to be the level set $\{x+iy\in\mathbbm{C}:G(x,y)=\log\rho\}$, where $G:\mathbbm{R}^{2}\to\mathbbm{R}$ is the unique solution to the exterior Laplace equation
$\nabla^{2}G=$  0 in $\,0\text{ in }\mathbbm{R}^{2}\setminus\Gamma,$
$G=$  0 on Γ, $\,0\text{ on }\partial\Gamma,$
$G(x)\sim$ $\,\log{\lvert x\rvert}\text{ as }{\lvert x\rvert}\to\infty.$ (36)
Furthermore, we let $E_{\rho}^{o}$ denote the open region bounded by $E_{\rho}$.
We note that, when $\Gamma=[a,b]\subset\mathbbm{R}$, the level set $E_{\rho}$ is a Bernstein ellipse with parameter $\rho$, with foci at $a$ and $b$. In Figure 6, we plot examples of level sets $E_{\rho}$ for an interval and for a sine curve, for various values of $\rho$.
(a) $\Gamma=[-1,1]$
(b) $\Gamma=\left\{t+0.3i\sin(3t):t\in[-1,1]\right\}$
Figure 6: The level set $E_{\rho}$ corresponding to $\Gamma$, for various values of $\rho$. The colorbar indicates the value of $\rho$. The smooth simple arc $\Gamma$ is the white curve in the figure. The plots were made using the source code provided in [2].
The following theorem illustrates just one situation where function approximation by polynomials over a smooth simple arc $\Gamma$ in the complex plane is feasible. We refer the readers to Section 4.5 in [28] for the proof.
Theorem 2.3.
Let $\Gamma$ be a smooth simple arc in the complex plane. Suppose that the function $F:\Gamma\to\mathbbm{C}$ is analytically continuable to the closure of the region $E_{\rho}^{o}$ corresponding to $\Gamma$, for some $\rho>1$. Then, there exists a sequence of polynomials $\{Q_{n}\}$ satisfying
${\lVert F-Q_{n}\rVert}_{L^{\infty}(\Gamma)}\leq C\rho^{-n},$ (37)
for all $n\geq 0$, where $C\geq 0$ is a constant that is independent of $N$.
@@
When $\Gamma$ is a line segment, the magnitude of the constant $C$ in (37) is proportional to ${\lVert F\rVert}_{L^{\infty}(E_{\rho}^{o})}$ (see Theorem 2.8 in Section 2.3). We conjecture that the same holds in the general case.
@@
Given a smooth simple arc $\Gamma\subset\mathbbm{C}$, define $\rho_{*}:=\inf\{\rho>1:D_{1}\subset E_{\rho}^{o}\}$, where $D_{1}$ is the open unit disk centered at the origin, and $E_{\rho}^{o}$ is the region corresponding to $\Gamma$ (see Definition 2.1).
The following lemma provides upper bounds for the 2-norm of the monomial coefficient vector of an arbitrary polynomial.
@@
Let $P_{N}:\mathbbm{C}\to\mathbbm{C}$ be a polynomial of degree $N$, where $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$ for some $a_{0},a_{1},\dots,a_{N}\in\mathbbm{C}$. The 2-norm of the coefficient vector $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$ satisfies
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert P_{N}\rVert}_{L^{\infty}(\partial D_{1})}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Gamma)},$ (38)
where $D_{1}$ denotes the open unit disk centered at the origin, and $\rho_{*}$ is given in Definition 2.2.
@@
where the last inequality comes from the fact that $D_{1}\subset E_{\rho_{*}}^{o}$ (see Definition 2.1). Finally, based on one of Bernstein’s inequalities (see Section 4.6 in [28]), we have
${\lVert P_{N}\rVert}_{L^{\infty}(E_{\rho_{*}}^{o})}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Gamma)}.$ (40)
@@
The following theorem provides an upper bound for the 2-norm of the monomial coefficients of an arbitrary interpolating polynomial.
@@
Let $\Gamma$ be a smooth simple arc in the complex plane, and let $F:\Gamma\to\mathbbm{C}$ be an arbitrary function. Suppose that there exists a finite sequence of polynomials $\{Q_{n}\}_{n=0,1,\dots,N}$, where $Q_{n}$ has degree $n$, which satisfies
${\lVert F-Q_{n}\rVert}_{L^{\infty}(\Gamma)}\leq C_{N}\rho^{-n},\quad 0\leq n\leq N,$ (41)
for some constants $\rho>1$ and $C_{N}\geq 0$. Define $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$ to be the $N$th degree interpolating polynomial of $F$ for a given set of distinct collocation points $Z=\{z_{j}\}_{j=0,1,\dots,N}\subset\Gamma$. The 2-norm of the monomial coefficient vector $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$ of $P_{N}$ satisfies
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert F\rVert}_{L^{\infty}(\Gamma)}+C_{N}\Bigl{(}\Lambda_{N}\Bigl{(}\frac{\rho_{*}}{\rho}\Bigr{)}^{N}+2\rho_{*}\sum_{j=0}^{N-1}\Bigl{(}\frac{\rho_{*}}{\rho}\Bigr{)}^{j}+1\Bigr{)},$ (42)
@@
Proof. Given $n\geq 0$, let $M^{(n)}:\mathbbm{R}^{n+1}\to\mathcal{P}_{n}$ be the bijective linear map associating each vector $(u_{0},u_{1},\dots,u_{n})^{T}\in\mathbbm{R}^{n+1}$ with the $n$th degree polynomial $\sum_{k=0}^{n}u_{k}z^{k}\in\mathcal{P}_{n}$. It follows immediately from Lemma 2.4 that, given any polynomial $P\in\mathcal{P}_{n}$,
${\bigl{\lVert}(M^{(n)})^{-1}[P]\bigr{\rVert}}_{2}\leq\rho_{*}^{n}{\lVert P\rVert}_{L^{\infty}(\Gamma)}.$ (43)
@@
=\,= ${\lVert Q_{0}\rVert}_{L^{\infty}(\Gamma)}+\sum_{j=0}^{N-1}{\bigl{\lVert}(M^{(j+1)})^{-1}[Q_{j+1}-Q_{j}]\bigr{\rVert}}_{2}$
@@
where the third inequality comes from the observation that $P_{N}-Q_{N}$ is the interpolating polynomial of $F-Q_{N}$ for the set of collocation points $Z$. $\blacksquare$
@@
The assumption (41) made in the theorem above can be satisfied for any function $F$ by choosing $C_{N}$ to be sufficiently large. When the function $F$ is continuable to the closure of the region $E_{\rho}^{o}$ corresponding to $\Gamma$ (see Definition 2.1), for some $\rho>\rho_{*}$, one can show that ${\lVert a^{(N)}\rVert}_{2}\leq 2{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}$, where $D_{1}$ is defined in Definition 2.2. This result comes from a generalization of Theorem 2.3 (see [10]), which says that ${\lVert F-P_{N}\rVert}_{L^{\infty}(E_{\rho_{*}}^{o})}=\mathcal{O}\bigl{(}(\rho_{*}/\rho)^{N}\bigr{)}$, from which it follows that
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert P_{N}\rVert}_{L^{\infty}(\partial D_{1})}\leq{\lVert P_{N}-F\rVert}_{L^{\infty}(\partial D_{1})}+{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}\leq 2{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}.$ (46)
The following theorem bounds the growth of the 2-norm of the inverse of a Vandermonde matrix.
@@
Suppose that $V^{(N)}\in\mathbbm{C}^{(N+1)\times(N+1)}$ is a Vandermonde matrix with $(N+1)$ distinct collocation points $Z=\{z_{j}\}_{j=0,1,\dots,N}\subset\mathbbm{C}$. Suppose further that $\Gamma\subset\mathbbm{C}$ is a smooth simple arc such that $Z\subset\Gamma$. The 2-norm of $(V^{(N)})^{-1}$ is bounded by
@@
where $\rho_{*}$ is given in Definition 2.2, and $\Lambda_{N}$ denotes the Lebesgue constant for the set of collocation points $Z$ over $\Gamma$.
@@
${\lVert a^{(N)}\rVert}_{2}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Gamma)}\leq\rho_{*}^{N}\Lambda_{N}{\lVert f^{(N)}\rVert}_{\infty}\leq\rho_{*}^{N}\Lambda_{N}{\lVert f^{(N)}\rVert}_{2},$ (48)
@@
Note that the bound above applies to any smooth simple arc $\Gamma\subset\mathbbm{C}$ that contains the set of collocation points $Z$.
Observation 2.4.
In the case where the set of $(N+1)$ collocation points $Z\subset\Gamma$ are chosen such that the associated Lebesgue constant $\Lambda_{N}$ is small, we observe in practice that the upper bound $\rho_{*}^{N}\Lambda_{N}$ is reasonably close to the value of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$ (see Figures 10 and 13(b) for numerical evidence).
@@
A worst case upper bound for ${\lVert a^{(N)}\rVert}_{2}$ is provided by inequality (48) in the proof of Theorem 2.6, as the right-hand side of this inequality is independent of the smoothness of the function $F$.
@@
Without loss of generality, we assume that the smooth simple arc $\Gamma$ is inside the unit disk centered at the origin (such that ${\lVert V^{(N)}\rVert}_{2}$ is small and $\gamma_{N}\lesssim 1$). Furthermore, we choose a set of $(N+1)$ collocation points $Z\subset\Gamma$ with a small Lebesgue constant $\Lambda_{N}$, and let $V^{(N)}$ denote the corresponding Vandermonde matrix. Recall from Theorem 2.2 that, if
@@
then the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}$ is bounded a priori by
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\lesssim{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+u\cdot{\lVert a^{(N)}\rVert}_{2},$ (51)
where $u$ denotes machine epsilon, $\widehat{P}_{N}$ is the computed monomial expansion, $P_{N}$ is the exact $N$th degree interpolating polynomial of $F$ for the set of collocation points $Z$, and $a^{(N)}$ is the monomial coefficient vector of $P_{N}$.
By Theorem 2.5, if there exists a constant $C_{N}\geq 0$ and a finite sequence of polynomials $\{Q_{n}\}_{n=0,1,\dots,N}$ such that ${\lVert F-Q_{n}\rVert}_{L^{\infty}(\Gamma)}\leq C_{N}\rho_{*}^{-n}$ for $0\leq n\leq N$, where $Q_{n}$ has degree $n$ and $\rho_{*}$ is given in Definition 2.2, then the monomial coefficient vector $a^{(N)}$ of $P_{N}$ satisfies
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Gamma)}\lesssim$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Gamma)}+u\cdot C_{N}.$ (53)

### v5

@@ 2 Polynomial interpolation in the monomial basis
As described in the introduction, when the Vandermonde system is solved using a backward stable linear system solver, the set of assumptions (11) and (12) is satisfied with a constant $\gamma_{N}\leq\zeta_{N}{\lVert V^{(N)}\rVert}_{2}$, where $\zeta_{N}>0$ is a growth factor bounded by a low-degree polynomial in $N$, and is independent of the matrix and the right-hand-side vector. It follows then that $\gamma_{N}\lesssim 1$. Therefore, given interpolation points with a Lebesgue constant $\Lambda_{N}\lesssim 1$, we have that ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\lesssim{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+u{\lVert a^{(N)}\rVert}_{2}$ when $\kappa(V^{(N)})\lesssim 1/u$, where we write the condition (14) as $\kappa(V^{(N)})={\lVert V^{(N)}\rVert}_{2}\cdot{\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{{\lVert V^{(N)}\rVert}_{2}}{2u\cdot\gamma_{N}}\lesssim 1/u$. In Figure 5, we plot the values of ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$, ${\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}$, and $u{\lVert a^{(N)}\rVert}_{2}$, for the functions appearing in Section 1, in order to validate the theorem above.
@@
(e) $F(x)={\lvert x+0.1\rvert}^{2.5}$
(f) $F(x)={\lvert\sin(5x)\rvert}^{3}$
Figure 5: Polynomial interpolation error, monomial approximation error and $u\cdot{\lVert a^{(N)}\rVert}_{2}$, for $\Omega=[-1,1]$. These functions are the ones that appear in Section 1.
@@
The second term on the right-hand side of (16) is an upper bound of the backward error ${\lVert P_{N}-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$, i.e., the extra loss of accuracy caused by the use of a monomial basis. Note that the absolute condition number of the evaluation of $P_{N}(z)$ in the monomial basis is $\lesssim{\lVert a^{(N)}\rVert}_{2}$ when ${\lvert z\rvert}\approx 1$, so that the resulting error is $\lesssim u\cdot{\lVert a^{(N)}\rVert}_{2}$, which is around the same size as $2u\cdot\gamma_{N}\Lambda_{N}{\lVert a^{(N)}\rVert}_{2}$.
Remark 2.2.
It is well-known that selecting a set of interpolation points with a small Lebesgue constant is crucial for the interpolating polynomial $P_{N}$ to approximate the function $F$ nearly as accurately as the optimal polynomial approximant [32]. This requirement on the Lebesgue constant $\Lambda_{N}$ of the interpolation points, that $\Lambda_{N}\lesssim 1$, can be circumvented by instead constructing an approximating polynomial using least-squares fitting. To construct such an approximating polynomial, one typically selects sample points on the boundary of the domain and uses a polynomial basis whose dimension is smaller than the number of sample points, as in, for example, [33]. Although our theoretical analysis only addresses approximations constructed using interpolation, the analysis for the least-squares case bears considerable similarity. The principal difference lies in the requirement for a more general notion of the Lebesgue constant. One such idea, using the 2-norm of the pseudoinverse mapping values at the sample points to the coefficients of an expansion in a basis, appears in [37].
In order to determine the relationship between the size of the monomial coefficient vector and the properties of the function being approximated, we will need the following definition, which generalizes the Bernstein ellipse to the case of a simply connected compact set in the complex plane.
@@
Given a simply connected compact set $\Omega$ in the complex plane and $\rho>1$, we define $E_{\rho}$ to be the level set $\{x+iy\in\mathbbm{C}:G(x,y)=\log\rho\}$, where $G:\mathbbm{R}^{2}\to\mathbbm{R}$ is the unique solution to the exterior Laplace equation
$\nabla^{2}G=$  0 in $\,0\text{ in }\mathbbm{R}^{2}\setminus\Omega,$
$G=$  0 on Ω, $\,0\text{ on }\partial\Omega,$
$G(x,y)\sim$ $\,\log r\text{ as }r:=\sqrt{x^{2}+y^{2}}\to\infty.$ (20)
We let $E_{\rho}^{o}$ denote the bounded open region with boundary $E_{\rho}$.
@@
[…]
When $\Omega=[a,b]\subset\mathbbm{R}$, the level set $E_{\rho}$ is a Bernstein ellipse with parameter $\rho$, with foci at $a$ and $b$. In Figure 6, we plot examples of level sets $E_{\rho}$ for various sets $\Omega$, for different values of $\rho$.
(a) $\Omega=\left\{t:t\in[-1,1]\right\}$
(b) $\Omega=\left\{re^{\pi it}:r\in[0,1],t\in[-1,1]\right\}$
(c) $\Omega=\left\{t+0.3i\sin(3t):t\in[-1,1]\right\}$
(d) $\Omega=\left\{r\bigl{(}\cos(\pi t)+0.2i\sin(\pi t)\bigr{)}:r\in[0,1],t\in[-1,1]\right\}$
(e) $\Omega=\left\{t+0.4i(t^{2}-1):t\in[-1,1]\right\}$
(f) $\Omega=\left\{a+bi:a,b\in[-\sqrt{2}/2,\sqrt{2}/2]\right\}$
Figure 6: The level sets $E_{\rho}$ for several values of $\rho$, corresponding to various sets $\Omega$. The level sets $E_{\rho}$ for $\rho=1,1.25,1.5,\dots$ appear as the boundaries between the colored regions in this figure. The color bar indicates which boundary corresponds to which value of $\rho$. The set $\Omega$ is shown in white. The dashed circle is the boundary of the open unit disk, used in the definition of $\rho_{*}$ in Definition 2.2. The plots were made using the source code provided in [2].
@@
Given a simply connected compact set $\Omega\subset\mathbbm{C}$, we define $\rho_{*}:=\inf\{\rho>1:D_{1}\subset E_{\rho}^{o}\}$, where $D_{1}$ is the open unit disk centered at the origin, and $E_{\rho}^{o}$ is the region corresponding to $\Omega$ defined in Definition 2.1.
The following lemma provides an upper bound for the 2-norm of the monomial coefficient vector of an arbitrary polynomial. We will use this lemma to prove a much tighter bound in Theorem 2.4.
@@
Let $P_{N}:\mathbbm{C}\to\mathbbm{C}$ be a polynomial of degree $N$, where $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$ for some $a_{0},a_{1},\dots,a_{N}\in\mathbbm{C}$. Let $D_{1}$ denote the open unit disk centered at the origin, and suppose that $\Omega\subset\mathbbm{C}$ is simply connected and compact. The 2-norm of the coefficient vector $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$ satisfies
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert P_{N}\rVert}_{L^{\infty}(\partial D_{1})}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Omega)},$ (21)
where $\rho_{*}$ is given in Definition 2.2.
@@
where the last inequality comes from the fact that $D_{1}\subset E_{\rho_{*}}^{o}$ (see Definition 2.1). Finally, using one of Bernstein’s inequalities (see Section 4.6 in [35]), we have that
${\lVert P_{N}\rVert}_{L^{\infty}(E_{\rho_{*}}^{o})}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Omega)}.$ (23)
@@
The following theorem is the principal result of this paper. It demonstrates that the upper bound for the 2-norm of the monomial coefficient vector of an arbitrary polynomial, provided in Lemma 2.3, can be significantly improved if the given polynomial is viewed as an interpolating polynomial of a function $F$. Instead of using only the size of ${\lVert P_{N}\rVert}_{L^{\infty}(\Omega)}$, this theorem uses the properties of the function $F$.
@@
Let $\Omega$ be a simply connected compact set in the complex plane, and let $F:\Omega\to\mathbbm{C}$ be an arbitrary function. Suppose that there exists a finite sequence of polynomials $\{Q_{n}\}_{n=0,1,\dots,N}$, where $Q_{n}$ has degree $n$, which satisfies
${\lVert F-Q_{n}\rVert}_{L^{\infty}(\Omega)}\leq C_{N}\rho^{-n},\quad 0\leq n\leq N,$ (24)
for some constants $\rho>1$ and $C_{N}\geq 0$. Define $P_{N}(z)=\sum_{k=0}^{N}a_{k}z^{k}$ to be the $N$th degree interpolating polynomial of $F$ for a given set of distinct interpolation points $Z=\{z_{j}\}_{j=0,1,\dots,N}\subset\Omega$. The 2-norm of the monomial coefficient vector $a^{(N)}:=(a_{0},a_{1},\dots,a_{N})^{T}$ of $P_{N}$ satisfies
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert F\rVert}_{L^{\infty}(\Omega)}+C_{N}\Bigl{(}\Lambda_{N}\Bigl{(}\frac{\rho_{*}}{\rho}\Bigr{)}^{N}+2\rho_{*}\sum_{j=0}^{N-1}\Bigl{(}\frac{\rho_{*}}{\rho}\Bigr{)}^{j}+1\Bigr{)},$ (25)
@@
Proof. Given $n\geq 0$, let $M^{(n)}:\mathbbm{C}^{n+1}\to\mathcal{P}_{n}$ be the bijective linear map associating each vector $(u_{0},u_{1},\dots,u_{n})^{T}\in\mathbbm{R}^{n+1}$ with the $n$th degree polynomial $\sum_{k=0}^{n}u_{k}z^{k}\in\mathcal{P}_{n}$. It follows immediately from Lemma 2.3 that, given any polynomial $P\in\mathcal{P}_{n}$,
${\bigl{\lVert}(M^{(n)})^{-1}[P]\bigr{\rVert}}_{2}\leq\rho_{*}^{n}{\lVert P\rVert}_{L^{\infty}(\Omega)}.$ (26)
@@
=\,= ${\lVert Q_{0}\rVert}_{L^{\infty}(\Omega)}+\sum_{j=0}^{N-1}{\bigl{\lVert}(M^{(j+1)})^{-1}[Q_{j+1}-Q_{j}]\bigr{\rVert}}_{2}$
@@
where the third inequality comes from the observation that $P_{N}-Q_{N}$ is the interpolating polynomial of $F-Q_{N}$ for the set of interpolation points $Z$. $\blacksquare$
@@
The assumption (24) made in Theorem 2.4 can be satisfied for any function $F$ by choosing $C_{N}$ to be sufficiently large. When the function $F$ has, for some $\rho>\rho_{*}$, an analytic continuation on the closure of the region $E_{\rho}^{o}$ corresponding to $\Omega$ (see Definitions 2.1 and 2.2), one can show that ${\lVert a^{(N)}\rVert}_{2}\lesssim 2{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}$, where $D_{1}$ is the open unit disk centered at the origin. This result is derived from Section 4.7 in [35], which establishes that ${\lVert F-P_{N}\rVert}_{L^{\infty}(E_{\rho_{*}}^{o})}=\mathcal{O}\bigl{(}(\rho_{*}/\rho)^{N}\bigr{)}$, where the asymptotic constant is proportional to ${\lVert F\rVert}_{L^{\infty}(E_{\rho_{*}}^{o})}$. Consequently, it follows that
${\lVert a^{(N)}\rVert}_{2}\leq{\lVert P_{N}\rVert}_{L^{\infty}(\partial D_{1})}\leq{\lVert F-P_{N}\rVert}_{L^{\infty}(\partial D_{1})}+{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}\lesssim 2{\lVert F\rVert}_{L^{\infty}(\partial D_{1})}.$ (29)
Theorem 2.4 provides an a priori upper bound on the 2-norm of the monomial coefficient vector. In order to apply Theorem 2.2, we also need to determine when the condition ${\lVert(V^{(N)})^{-1}\rVert}_{2}\leq\frac{1}{2u\cdot\gamma_{N}}$ applies. The following theorem, which is an immediate corollary of Lemma 2.3, bounds the growth of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$. It includes some previous results [14, 15] as special cases of our bound.
@@
Suppose that $V^{(N)}\in\mathbbm{C}^{(N+1)\times(N+1)}$ is a Vandermonde matrix with $(N+1)$ distinct interpolation points $Z=\{z_{j}\}_{j=0,1,\dots,N}\subset\mathbbm{C}$. Suppose further that $\Omega\subset\mathbbm{C}$ is a simply connected compact set such that $Z\subset\Omega$. The 2-norm of $(V^{(N)})^{-1}$ is bounded by
@@
where $\rho_{*}$ is given in Definition 2.2, and $\Lambda_{N}$ denotes the Lebesgue constant for the set of interpolation points $Z$ over $\Omega$.
@@
${\lVert a^{(N)}\rVert}_{2}\leq\rho_{*}^{N}{\lVert P_{N}\rVert}_{L^{\infty}(\Omega)}\leq\rho_{*}^{N}\Lambda_{N}{\lVert f^{(N)}\rVert}_{\infty}\leq\rho_{*}^{N}\Lambda_{N}{\lVert f^{(N)}\rVert}_{2},$ (31)
@@
Note that the bound above applies to any simply connected compact domain $\Omega\subset\mathbbm{C}$ that contains the set of interpolation points $Z$.
Observation 2.5.
In the case where the set of $(N+1)$ interpolation points $Z\subset\Omega$ are chosen such that the associated Lebesgue constant $\Lambda_{N}\lesssim 1$, we observe in practice that the upper bound $\rho_{*}^{N}\Lambda_{N}$ is close to the value of ${\lVert(V^{(N)})^{-1}\rVert}_{2}$ (see Figures 10 and 14 for numerical evidence).
@@
A worst case upper bound for ${\lVert a^{(N)}\rVert}_{2}$ is provided by inequality (31) in the proof of Theorem 2.5, as the right-hand side of this inequality depends only on the size of $F$.
Remark 2.7.
Formula (4.5) in [7] establishes the bound ${\lVert(V^{(N)})^{-1}\rVert}_{\infty}\leq\Delta_{N}(D_{1},\Omega)\cdot\Lambda_{N}$, where $\Delta_{N}(D_{1},\Omega):=\sup_{p\in\mathcal{P}_{N},p\neq 0}\frac{{\lVert p\rVert}_{L^{\infty}(D_{1})}}{{\lVert p\rVert}_{L^{\infty}(\Omega)}}$, and $\mathcal{P}_{N}$ denotes the set of all polynomials of degree less than or equal to $N$. By combining the maximum principle with Lemma 2.3, we obtain the bound $\Delta_{N}(D_{1},\Omega)\leq\rho_{*}^{N}$, from which it follows that ${\lVert(V^{(N)})^{-1}\rVert}_{\infty}\leq\rho_{*}^{N}\Lambda_{N}$. This result can alternatively be obtained more directly using the inequality
${\lVert(V^{(N)})^{-1}\rVert}_{\infty}=\sup_{f^{(N)}\neq 0}\Bigl{\{}\frac{{\lVert a^{(N)}\rVert}_{\infty}}{{\lVert f^{(N)}\rVert}_{\infty}}\Bigr{\}}\leq\sup_{f^{(N)}\neq 0}\Bigl{\{}\frac{{\lVert a^{(N)}\rVert}_{2}}{{\lVert f^{(N)}\rVert}_{\infty}}\Bigr{\}}\leq\rho_{*}^{N}\Lambda_{N},$ (33)
where the final inequality follows from (31).
@@
As before, we assume that the simply connected compact set $\Omega$ is inside the unit disk centered at the origin (such that $\gamma_{N}\lesssim 1$). Furthermore, we choose a set of $(N+1)$ interpolation points $Z\subset\Omega$ with a Lebesgue constant $\Lambda_{N}\lesssim 1$, and let $V^{(N)}$ denote the corresponding Vandermonde matrix. Recall from Theorem 2.2 that, if
@@
then the monomial approximation error ${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}$ is bounded a priori by
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\lesssim{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+u\cdot{\lVert a^{(N)}\rVert}_{2},$ (35)
where $u$ denotes machine epsilon, $\widehat{P}_{N}$ is the computed monomial expansion, $P_{N}$ is the exact $N$th degree interpolating polynomial of $F$ for the set of interpolation points $Z$, and $a^{(N)}$ is the monomial coefficient vector of $P_{N}$.
By Theorem 2.4, if we choose a constant $C_{N}\geq 0$ and a finite sequence of polynomials $\{Q_{n}\}_{n=0,1,\dots,N}$ such that ${\lVert F-Q_{n}\rVert}_{L^{\infty}(\Omega)}\leq C_{N}\rho_{*}^{-n}$ for $0\leq n\leq N$, where $Q_{n}$ has degree $n$ and $\rho_{*}$ is given in Definition 2.2, then the monomial coefficient vector $a^{(N)}$ of $P_{N}$ satisfies
@@
${\lVert F-\widehat{P}_{N}\rVert}_{L^{\infty}(\Omega)}\lesssim$ $\,{\lVert F-P_{N}\rVert}_{L^{\infty}(\Omega)}+u\cdot C_{N}.$ (37)